# Manify: A Python Library for Learning Non-Euclidean Representations

### v1

@@ G.1.1 Validating Coordinate Learning (Gu et al 2018)
G.2 Validating $\kappa$-GCN Model Relations
In Table 5, we test the relationships laid out in Figure 3. We applied the Wilcoxon rank-sum test over prediction accuracies for Gaussian mixtures with 3 clusters and 2 classes in $\mathbb{E}^{2}$ for 30 trials apiece. We define the models as follows:
  • The $\kappa$-MLP is a $\kappa$-GCN with $\kappa=0$ and $\mathbf{\hat{A}}=\mathbf{I}$, and one hidden layer;
@@ G.2 Validating $\kappa$-GCN Model Relations
  • The Pytorch MLP is a Pytorch neural network with one hidden layer. To match $\kappa$-GCN, it has square weights and no biases;
@@
  • The “Scikit-Learn MLR” refers to the ‘LogisticRegression’ class in Scikit-Learn.
All predictors were trained for 2,000 epochs.
Table 5: Wilcoxon rank-sum test $p$-values for accuracies of different models across 30 trials of Gaussian mixture classification. By default, we expect MLP models to be similar to one another, and MLR models to be similar to one another, but MLR and MLP models to differ.
Pytorch Pytorch Scikit-Learn
$\kappa$-MLP MLP $\kappa$-MLR MLR MLR
$\kappa$-MLP 1.00 0.477 1.00 0.477 0.142
Torch MLP 0.477 1.00 0.477 0.477 0.524
$\kappa$-MLR 1.00 0.477 1.00 1.00 0.142
Torch MLR 0.477 0.477 1.00 1.00 0.142
Scikit-Learn MLR 0.142 0.524 0.142 0.142 1.00

### v2

@@ H.2 Validating Coordinate Learning (Gu et al 2018)
H.3 Validating $\kappa$-GCN Model Relations
In Table 5, we test the relationships laid out in Figure 3. We applied the Wilcoxon rank-sum test over prediction accuracies for Gaussian mixtures with 3 classes in $\mathbb{E}^{2}$ for 10 trials apiece. We define the models as follows:
  • The $\kappa$-GCN uses a radial basis function rescaled by 0.1 for $\hat{\mathbf{A}}$. This scaling factor was chosen to ensure $\hat{\mathbf{A}}$ does not resemble $\mathbf{I}$, preventing accidental resemblance to MLPs. We use four $\kappa$-GCN layers, followed by a stereographic logits layer;
  • The $\kappa$-MLP is a $\kappa$-GCN with $\kappa=0$ and $\mathbf{\hat{A}}=\mathbf{I}$, and four hidden layers;
@@ H.3 Validating $\kappa$-GCN Model Relations
  • The Pytorch GCN implements simple graph convolution layers with an analogous architecture and $\hat{\mathbf{A}}$ to $\kappa$-GCN. To match $\kappa$-GCN, it has square weights and no biases;
  • The Pytorch MLP is a Pytorch neural network with four hidden layers. To match $\kappa$-GCN, it has square weights and no biases;
@@
  • The “Scikit-Learn MLR” refers to the ‘LogisticRegression’ class in Scikit-Learnwith no regularization.
All predictors were trained for 1,000 epochs with a learning rate of 0.01 (Pytorch models) and 0.025 ($\kappa$-GCN derived models). We use 0.025 to counteract the fact that $\kappa$-GCN rescales logits by a factor of 4, making the effective learning rates equivalent.
Table 5: Wilcoxon rank-sum test $p$-values for accuracies of different models across 10 trials of Gaussian mixture classification. By default, we expect all GCN, MLP, and MLR models to be similar to one another (high $p$-value, and dissimilar to all other models. We color $p$-values less than 0.1 in red, and $p$-values greater than 0.9 in green. The observed green clusters reflect our expectations, except that Scikit-Learn MLR diverges from $\kappa$- and Pytorch MLR.
$\kappa$-GCN Pytorch $\kappa$-MLP Pytorch $\kappa$-MLR Pytorch Scikit-Learn
GCN MLP MLR MLR
$\kappa$-GCN 1 1 0.0625 0.375 0.0098 0.0098 0.0098
Pytorch GCN 1 1 0.7422 0.125 0.0059 0.0059 0.0059
$\kappa$-MLP 0.0625 0.7422 1 0.9453 0.082 0.082 0.084
Pytorch MLP 0.375 0.125 0.9453 1 0.0488 0.0488 0.0488
$\kappa$-MLR 0.0098 0.0059 0.082 0.0488 1 1 0.5547
Pytorch MLR 0.0098 0.0059 0.082 0.0488 1 1 0.625
Scikit-Learn MLR 0.0098 0.0059 0.084 0.0488 0.5547 0.625 1